Determine what type of number the solutions to the equation 4x^2+7x=0 are, and how many (distinct) solutions exist.
Answer by checkley77(12844)

4x^2+7x=0
x(4x+7)=0
x=0 ans.
4x+7=0
4x=-7
x=-7/4 ans.